Two numbers add up to 48. One of the numbers is -22, find the other.
Answer by Alan3354(69443)

Two numbers add up to 48. One of the numbers is -22, find the other.
-------------------
-22 + 0 = -22 NG
-22 + 22 = 0 NG
-22 + 44 = 22 NG but getting closer
---
You try some.